OpenStudy (anonymous):

Solve the system using the substitution method. y-3x=2 and y=(x+1)^2-5

OpenStudy (anonymous):

So y is already isolated for you in the second equation, so substitute (x+1)^2-5 into the first equation

OpenStudy (anonymous):

\[y-3x=2 \\ y=(x+1)^2-5\] right?

OpenStudy (anonymous):

first equation solve for \(y\) get \[y=3x+2\]

OpenStudy (anonymous):

^2-5

OpenStudy (anonymous):

next step is to set them equal and write \[(x+1)^2-5=3x+2\] which is a quadratic equation do you know how to solve it?

OpenStudy (anonymous):

thank you so much!!